OpenStudy (lolacole12):

What is the value of the expression –9 – 7 + 5? A. 3 B. 11 C. –11 D. –7

OpenStudy (lolacole12):

@mathstudent55

OpenStudy (lolacole12):

I think its d right

OpenStudy (lolacole12):

o no its c

OpenStudy (mathstudent55):

You need to know the order of operations. Have you learned order of operations? Many people remember it as PEMDAS. You have only subtractions and additions. Do the problem in the order the operations occur from left to right.

OpenStudy (mathstudent55):

-9 - 7 + 5 = -16 + 5 = -11 You are correct.

OpenStudy (mathstudent55):

C is correct.

OpenStudy (lolacole12):

thanks one more

OpenStudy (mathstudent55):

ok